OpenStudy (anonymous):

Algebra 1 multiple choice word problem.

OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8 A. −8 + y B. −8 − y C. 8 + y D. 8 − y

OpenStudy (anonymous):

−x − y = −8 Each side of the above equation multiplied by -1: x + y = 8 Can you take it from here?

OpenStudy (anonymous):

D. perhaps?